Quaker State Wings has 220,000 shares outstanding and net income of $880,000. The company stock is currently selling for $85.57 per share. If the company repurchases $614,000 of its stock, what is the earnings per share after the repurchase?
Answer #1

Stock repurchased = amount/price = 614000/85.57 = 7175.41 = 7175 shares

EPS = Net Income/Shares outstanding

= 880000/(220000-7175)

= $4.13